OpenStudy (anonymous):

Max is 5 years older than Paulette. Next year he will be twice as old as she will be. How old is each now?

OpenStudy (anonymous):

Well next year he will be one year older than he is now he will still be 5 years older and twice the age of paulette. So think of the number where adding 5 is the same as doubling and then remove one year from your answer to get paulette's age. And add 5 to get max's age

OpenStudy (anonymous):

how do i set up the equation though?

OpenStudy (anonymous):

the way i set it up is: m=5p and for the second equation i put: m+1=2(p+1)

OpenStudy (anonymous):

consider x to be paulettes age now and then next year x+1 then maxs age now is x+5 and then next year will be 2(x+1) solve as a system

OpenStudy (anonymous):

m = p+5 m+1 = 2(p+1) and solve the simultaneous eqns so taking eqn 1 from eqn 2 gives 1 = p-3 and hence p=4 and substituting back into eqn 1 gives m=4+5=9

OpenStudy (anonymous):

oh i thought that the first equation was 5p because it was five years older so i thought it was multipilying

OpenStudy (anonymous):

word problems can always be a little tricky

OpenStudy (anonymous):

yes and the thing that bothered me was that i was not at scholl the day i was suposed to learn this

OpenStudy (anonymous):

yeah it's always better to have explainations before you dive into homework. if you have any more problems that you need help with that are similar i'd be happy to help :)

OpenStudy (anonymous):

ok thank you